OpenStudy (anonymous):

PDE's: Show that v(x,y) = xF(2x+y) is a general solution of xv_x - 2xv_y = v, where F is an arbitrary differentiable function.

OpenStudy (anonymous):

Attached is my solution, somehow I did it wrong, please point my mistake out.

OpenStudy (jamesj):

I'm pretty sure there's an error in your PDE. It should be \[ xv_x - 2x^2v_y = v \] This is a change of -2x in front of v_y to -2x^2 Now, \[ v(x,y) = xF(2x+y) \] \[ v_x = F(2x + y) + 2xF'(2x + y) \] \[ v_y = F'(2x + y) \] Write F for F(2x + y), F' for F'(2x + y) Hence \[ xv_x - 2x^2v_y \] \[ = xF + 2x^2F' - 2x^2F' \] \[= xF\] \[ = v\]

OpenStudy (anonymous):

Hm, thanks, I'll check if there's an errata list on this book, I have a very old copy (though same ISBN as the newer copies) of Schaum's Outline Series for Fourier Analysis with applications to BVPs.

OpenStudy (anonymous):

But really, I copied the exact equation in the book, except for the notation.

OpenStudy (jamesj):

I believe you. But as it is, that form of v does not satisfy the equation.

OpenStudy (anonymous):

Haha thank you, I'll tell my prof.

OpenStudy (jamesj):

Wait, wait. I made a mistake. \[ v_y = xV'(2x+y) \]

OpenStudy (jamesj):

Now, substitute that back in and it DOES satisfy the equation. Apologies.

OpenStudy (jamesj):

For example, let F(t) = t, and thus \[ v = xF(2x+y) = x(2x+y) = 2x^2 + xy \] Hence \[ v_x = 4x + y, v_y = x \] Therefore \[ xv_x - 2xv_y = x(4x + y) - 2x^2 \] \[ = 2x^2 + xy \] \[ = x(2x + y) = xF(2x+y) \]

OpenStudy (anonymous):

Haha thanks, my prof also pointed out that the way I was doing it was making the problem much more difficult than it really is. Now my problem is finding a particular solution if v(1,y)=y^2.